What is the present value of the following future amounts? 

a. $700 to be received 8 years from now discounted back to the present at 10 percent 

b. 5400 to be received 7 years from now discounted back to the present at 7 percent 

c. $1,200 to be received 12 years from now discounted back to the present at 4 percent 

d. $1,200 to be received 5 years from now discounted back to the present at 18 percent 


a. What is the present value of $700 to be received 8 years from now discounted back to the present at 10 percent? (Round to the nearest cent.) 

b. What is the present value of $400 to be received 7 years from now discounted back to the present at 7 percent? (Round to the nearest cent) 

c. What is the present value of $1 200 to be received 12 years from now discounted back to the present at 4 percent? (Round to the nearest cent) 

d. What is the present value of $1.200 to be received 5 years from now discounted back to the present at 18 percent? (Round to the nearest cent)
Answer #1

Answer $ 700 a) amount to be recieved in each year 10% annuity factor for 8 years including this year It (1+0.1)? On 5.8684 nc) amount to be recieved in each yecy = $1200 4% annuity factor for 12 years from now 1 1+ (1+0.04) = 9.7605 Present value =
Answer #2

SOLUTION :


a.


Discount rate, r = 10% = 0.10 


PV 

= FV / (1 + r)^n

= 700 / ( 1 + 0.10)^8

= 326.56 ($) 


So, PV of $700 to be received 8 years from now and discounted back to the present at the rate of 10% is = 326.56 ($) (ANSWER)


b.


Discount rate, r = 7% = 0.07 


PV 

= FV / (1 + r)^n

= 5400 / ( 1 + 0.07)^7

= 3362.85 ($) 


So, PV of $5400 to be received 7 years from now and discounted back to the present at the rate of 7% is = 3362.85 ($) (ANSWER)


c.


Discount rate, r = 4% = 0.04 


PV 

= FV / (1 + r)^n

= 1200 / ( 1 + 0.04)^12

= 749.52 ($) 


So, PV of $1200 to be received 12 years from now and discounted back to the present at the rate of 4% is = 749.52 ($) (ANSWER).


d.


Discount rate, r = 5% = 0.05 


PV 

= FV / (1 + r)^n

= 1200 / ( 1 + 0.05)^18

= 498.62 ($) 


So, PV of $1200 to be received 18 years from now and discounted back to the present at the rate of 5% is = 498.62 ($) (ANSWER).




answered by: Tulsiram Garg